write an equation for a line which has y-intercept -6 and x-intercept 2
Answer by oberobic(2304)

y = mx + b
.
b = -6
.
y = mx -6
.
The line goes through (0,-6) and (2,0)
.
m = -6-0 /0-2
m = 3
.
y = 3x -6
.
Done.